# bibliography.bib

### v1

@@ Abstract
Mathematics Subject Classifications. 65Y20, 65C05, 65C30, 65G20, 60-08
@@ 3.2 Approximate FE solutions using inexact solvers
Due to the limitations of inexact linear solvers, the discrete solution $u_{h}$ is not obtained exactly in practice. Instead, we compute an approximation $\widehat{u}_{h}$. The aim of this section is to estimate the error $\smash{\|u_{h}(\cdot,\omega)-\widehat{u}_{h}(\cdot,\omega)\|_{H_{0}^{1}(D)}}$ by means of the residual of the solution of the FE system $Ax=b$ and investigate how iterative refinement from Section 2.3 can be employed to solve this system. Let us first introduce some notation.
@@ 4.1 Standard multilevel Monte Carlo
Suppose $Q:\Omega\rightarrow\mathbb{R}$ is a random variable and we are interested in computing its mean $\mathbb{E}[Q]$. Assume that $Q$ cannot be evaluated sample-wise, but a sequence of models $Q_{l},l\in\{0,\ldots,N\}$, approximating $Q$ with increasing accuracy, are available. We define the following auxiliary Monte Carlo (MC) estimators:
@@
will then be referred to as the MLMC estimator for $E[Q]$. Although not necessary, the estimators $\widehat{Y}_{l}$ are, in this work, assumed to be independent. In this definition, it is assumed that the models can be evaluated exactly, i.e., we can compute $Q_{l}(\omega)$.
@@
It can be shown that three complexity regimes can be distinguished based on the values of $\beta$ and $\gamma$. The optimal performance of the MLMC algorithm is achieved when $\beta>\gamma$, i.e., when variance decays faster than cost increases.
The full complexity analysis can be found in [cliffe2011multilevel, Theorem 1].
@@ 4.3 Application to the elliptic PDE problem
Let $G:H_{0}^{1}(D)\rightarrow\mathbb{R}$ be a bounded linear functional and consider Problem 3.1, an AVP with random data and solution $u(\cdot,\omega)\in H_{0}^{1}(D)$ for a.e. $\omega\in\Omega$. We consider the problem of estimating the quantity of interest (QoI) defined as the expected value of the random variable $Q:\Omega\rightarrow\mathbb{R}$ given by $\smash{\omega\mapsto G(u(\cdot,\omega))}$.
@@
Let $m>1$, and let $h_{0},h_{1},\ldots$ be discretisation parameters satisfying $h_{0}>0$ and $\smash{mh_{l}=h_{l-1}}$. Let $\widehat{u}_{h_{l}}$ be the discrete solution of the AVP with random data (Problem 3.1) computed effectively to precision $\delta_{l}$ on mesh level $l$, and assume that there are $k_{1},k_{2}>1$ such that $k_{1}\delta_{l}\leq\delta_{l-1}\leq k_{2}\delta_{l}$ for all $l\geq 1$. Then
@@ 4.4 Adaptive MPML algorithm
for a fixed constant $k_{p}\in(0,1)$. To balance the terms in the error estimate, it is sufficient to choose $\smash{\delta_{l},\;l=0,\ldots,L}$ such that
@@
0:  $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$
@@
7:     Update estimates for $N_{l}$ as $N_{l}\coloneqq\sqrt{\frac{V_{l}}{C_{l}}}\frac{2}{TOL^{2}}\sum_{k=0}^{L}\sqrt{V_{k}C_{k}}$
8:     if $|\widehat{Y}_{L}|>\frac{rm^{\alpha}-1}{\sqrt{2}}TOL$ then
@@
12:     if $|\widehat{Y}_{L}|\leq\frac{rm^{\alpha}-1}{\sqrt{2}}TOL$ and $\sum_{l=0}^{L}s^{2}_{l}/N_{l}\leq TOL^{2}/2$ then
@@ 5 Numerical results
We provide numerical experiments to demonstrate the potential cost savings using the adaptive MPML algorithm (Algorithm 4.1) compared to the standard adaptive MLMC algorithm while preserving the overall computational accuracy. To this end we use a model given by an elliptic PDE with lognormal random coefficients with both a direct and an iterative linear solver. All numerical experiments are implemented in Python. Our Python codes are available at https://github.com/josef-martinek/mpml.
@@ 00 $hhss$
Table 1: Left: Required effective precision (in terms of relative residual) on each level determined by the adaptive MPML algorithm (Algorithm 4.1) for different values of the finest level $L$.
Right: Choice of the precisions for iterative refinement. Note that the factorisation is carried out in half precision on the coarsest level.
Due to the fact that $\omega_{j}\sim N(0,\sigma^{2})$, assumptions (4) and (5) are not satisfied. One could choose, for example, $\omega_{j}\sim\text{Uni}(0,c)$, in which case both assumptions are satisfied. However, we want to test the developed methods under more realistic circumstances.
We will be concerned with the quantity of interest given by
@@
A thorough analysis of MLMC applied to a PDE of the form (27) has been presented in [teckentrup2013further]. For a rigorous analysis, it is further necessary to modify the results from Section 3.2, which is not presented here.
@@
In the first example of this section, we choose the data in (27) as follows. The right-hand side satisfies $f\equiv 1$ on $D$ and the parameters in the coefficient function are chosen as $s=4$, $q=2$, and $\sigma=2$. To solve this problem numerically, we employ the FEM for discretisation as described in Section 3 with simplical elements implemented in FEniCSx [dolfinx_2023]. In this experiment the initial discretisation parameter is $h_{0}=1/8$ and the mesh is refined on each level by a factor $m=2$.
@@
Recall that throughout our experiments by standard MLMC we mean Algorithm 4.1 without step 2, where the computations in step 4 are carried out with an a-priori given high accuracy. In each experiment we always describe precisely what the accuracy is. The parameters $\alpha$ and $\beta$ in MLMC from the MLMC complexity theorem (see (10) and [cliffe2011multilevel, Theorem 1]) can be chosen as $\beta=4$, $\alpha=2$ for the random field in (28). To set up the MPML algorithm (Algorithm 4.1), we further need the parameters $\alpha_{1}$$\alpha_{2}$$\beta_{1}$, and $\beta_{2}$ from the MPML complexity theorem (Theorem 4.1). As shown in Lemma 4.3, we have $\alpha_{1}=\alpha=2$, $\beta_{1}=\beta=4$, $\alpha_{2}=1$, and $\beta_{2}=2$. The cost of solving the linear system is in both MLMC and MPML adaptive algorithms estimated as the number of floating point operations (FLOPs) performed by MINRES. To this end we use the PETSc GetFlops() function. The underlying linear system is solved using the PETSc implementation of MINRES (see [petsc-web-page], [petsc_4_py]) with a stopping criterion given by the relative residual norm. The values of the stopping criterion in the MPML algorithm are given in Table 1, left. The standard MLMC algorithm uses a fixed value specified below for each experiment. Let us point out that in this case all computations are carried out in double precision and no iterative refinement is used.
@@
Figure 1: MPML MSE compared to the reference MLMC MSE and the given MSE tolerance. Both method use MINRES as linear solver, MPML uses adaptively chosen stopping criterion. Approximate $95\%$ confidence intervals are shown.
@@
Figure 2: Total cost gain in terms of FLOPs for various tolerances using adaptive MPML compared to the standard adaptive MLMC. Both methods use MINRES as linear solver, MPML uses adaptively chosen stopping criterion. Approximate $95\%$ confidence intervals are shown.
@@
To determine the effective precision $\delta_{l}$ in the MPML adaptive algorithm, we use the formula (24). To this end we need to specify the constant $k_{p}$. In this experiment we set $k_{p}=0.05$. As discussed below and depicted in Figure 4, the MPML algorithm is not sensitive with respect to the choice of $k_{p}$. For a multilevel estimator with a given number of levels, the choice $k_{p}=0.05$ now determines uniquely the effective precisions $\delta_{l}$ on all levels according to (24). The values of $\delta_{l}$ for different estimators can be found in Table 1.
@@
We now aim to verify the accuracy of the MPML algorithm. We execute $1000$ runs of the adaptive MPML algorithm (Algorithm 4.1) for each of the four values of the MSE tolerance $TOL^{2}$, namely $\smash{TOL^{2}=8,4,2,1\times 10^{-6}}$. For the same tolerances we perform $1000$ runs of the standard MLMC algorithm. The standard MLMC algorithm uses a fixed value $10^{-6}$ for the MINRES stopping criterion. Figure 1 shows the MSE of the estimates obtained by both algorithms averaged over the number of runs with approximate $95\%$ confidence intervals. To estimate the MSE of both algorithms a reference value of the QOI is computed by the standard MLMC algorithm with the tolerance $TOL^{2}=2\times 10^{-8}$ with PETSc LU factorisation as the linear solver. We observe that the error of both MPML and MLMC algorithm is close to the tolerance within statistical error.
@@
In the following we measure the cost gains obtained by MPML (Algorithm 4.1) compared to standard MLMC. We plot the average cost in terms of FLOPs of the $1000$ final estimators obtained for each tolerance by the adaptive algorithms against the respective MSE tolerances. In Figure 2 we observe that the cost is reduced by a factor of $\approx 1.5$. As demonstrated in Figure 1, this cost gain comes with no statistically significant loss of accuracy. The cost gain comes from the required effective precision $\delta_{l}$ of solving the linear system (in terms of relative residual) being much larger than a typically chosen convergence criterion, see Table 1. As above, the standard MLMC algorithm uses a fixed value $10^{-6}$ of the MINRES stopping criterion.
@@
At the beginning of the cost analysis in Section 4.5 we made the assumption that the MPML and the MLMC adaptive algorithms produce estimators with approximately the same number of samples on each level. We verify numerically that this is true. Figure 3 shows the average number of samples obtained by MPML (Algorithm 4.1) and standard MLMC on each level for the tolerance $TOL^{2}=1\times 10^{-6}$. We observe that the number of samples are equal within statistical error given by the approximate $95\%$ confidence intervals. For this experiment, the standard MLMC algorithm uses a fixed value $10^{-10}$ of the MINRES stopping criterion.
@@
Figure 3: Average number of samples MLMC and MPML algorithms converge to for the MSE tolerance $10^{-6}$. Approximate $95\%$ confidence intervals are shown.
@@
Figure 4: MSE of the MPML algorithm for various values of the constant $k_{p}$ compared to the reference MLMC MSE. Approximate $95\%$ confidence intervals are shown.
@@
In Section 4.4 we claimed that the MPML algorithm is not sensitive with respect to the a-priori chosen constant $k_{p}$ needed to determine the required computational accuracy. Figure 4 shows the average MSE of the MPML algorithm (Algorithm 4.1) for the MSE tolerance $TOL^{2}=2\times 10^{-6}$. The MSE is estimated using $1000$ runs of the algorithm for the values $k_{p}=0.05,0.1,0.2,0.4$. The estimated MSE of the standard MLMC algorithm is shown for comparison, with approximate $95\%$ confidence intervals. The reference MLMC MSE is computed using $1000$ runs of standard MLMC algorithm using MINRES requiring that the relative residual is smaller than $10^{-10}$. It can be concluded that the choice of the constant $k_{p}$ has a minimal impact on the overall accuracy as long as the constant is sufficiently small. Our suggested choice $k_{p}=0.05$ bounds the computational error in our model problem safely.
5.2 Elliptic PDE with Lognormal Random Coefficients – Direct Solver
@@
In this section we again solve the problem (27) using the same input data, i.e., the right-hand side satisfies $f\equiv 1$ on $D$ and the parameters in the coefficient function are chosen as $s=4$, $q=2$, and $\sigma=2$. The initial discretisation parameter is again chosen as $h_{0}=1/8$. However, here the underlying linear system (with an SPD matrix) is solved using a double precision Cholesky factorisation from PETSc in the standard MLMC algorithm and a low precision Cholesky factorisation with iterative refinement in our MPML algorithm. Recall that throughout our experiments by standard MLMC we mean Algorithm 4.1 without step 2, where the computations in step 4 are carried out with an a-priori given high accuracy. The implementation of iterative refinement and the Cholesky factorisation used within it is our own implementation. To carry out computations in half, single, and double precision, we use the numerical types of NumPy. Namely, we use float16(), float32(), and float64() for half, single, and double precision, respectively. Apart from the linear solver, all calculations are carried out in double precision.
@@
We set up the iterative refinement (Algorithm 2.1) as follows. The numerical value of the stopping criterion is, analogously to the previous section, given by the required effective precision $\delta_{l}$ specified by formula (24). To this end, we choose $\alpha_{1}=\alpha=2$, $\beta_{1}=\beta=4$, $\alpha_{2}=1$, and $\beta_{2}=2$ as in the previous example. Table 1 (left) shows the resulting values of the effective precision $\delta_{l}$ used in the iterative refinement algorithm as the stopping criterion. Table 1 (right) shows the exact setting of the iterative refinement. The algorithm contains $3$ precisions, i.e., $\varepsilon_{f}$, $\varepsilon$, and $\varepsilon_{r}$ that take on one of the values quarter ($q$), half ($h$), single ($s$), or double ($d$). To simplify the notation, we describe the exact setting of the iterative refinement schematically as an ordered quadruple, e.g., $(\varepsilon_{f},\varepsilon_{7},\varepsilon,\varepsilon_{r})=(hhss)$, where $\varepsilon_{7}$ is the precision chosen at step 8 of Algorithm 2.1.
@@
Figure 5: MPML MSE compared to the reference MLMC MSE using the same number of samples and the same random seed. MLMC uses double precision Cholesky, MPML uses low precision Cholesky with iterative refinement. Approximate $95\%$ confidence intervals are shown.
@@
Figure 6: Cholesky + Iterative refinement MPML and double precision Cholesky MLMC memory access cost comparison. Total cost of the estimator per level is shown (cost of $1$ sample $\times$ number of samples).
@@
We now want to determine accuracy of the MPML estimator from Algorithm 4.1. Here we aim to eliminate randomness from estimating the MSE by fixing a random seed to get a better idea of how big the computational error actually is. The procedure is as follows. We perform $1000$ runs of the standard MLMC algorithm using the cost model $\gamma=2$. This model corresponds to the case when the linear system is solved in linear complexity. The algorithm is run for each of the values of the MSE tolerance $TOL^{2}$, namely $TOL^{2}=8,4,2,1\times 10^{-6}$. For each tolerance, we then average over the resulting $1000$ estimators to obtain a single multilevel estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ with a fixed (average) number of samples. We then perform $1000$ runs of the $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ estimator using double precision Cholesky as the linear solver. We also execute $1000$ runs of the corresponding $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ estimator with the same number of samples as the $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ estimator and the same random seed. Figure 5 shows the resulting average mean squared errors of both estimators. We conclude that the difference in the MSEs of both estimators is small; in fact the relative error of the MPML MSE with respect to the MLMC MSE is less than $0.01\%$. Recall that the MPML estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ uses half and single precision for the matrix factorisation on all levels; see Table 1 (right). This promises significant memory savings when implemented efficiently on an architecture where half precision computations are supported.
@@
We will now estimate the cost gains using MPML (Algorithm 4.1) with low precision Cholesky factorisation and iterative refinement compared to standard MLMC with double precision Cholesky. The gains are estimated using a very simple cost model, assuming that the cost of a memory access is directly proportional to the number of bits used to store the number in memory. In this model, for example, accessing a half precision floating point number is 2×2$\times$2 × cheaper than accessing a single precision number and 4×4$\times$4 × cheaper than accessing a double precision number. We can then simulate the cost gain by counting the non-zeros computed and stored in the process of solving the linear system $Ax=b$. Figure 6 shows the total cost gain per level (cost per sample $\times$ number of samples) of the $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ compared to the standard $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$. On all levels we observe a simulated memory gain of $\approx 2$. The total memory gain (sum over all levels) is $\approx 2$ as well.
@@
Figure 7: MPML MSE compared to the reference MLMC MSE using the same number of samples and the same random seed. MLMC uses double precision Cholesky, MPML uses low precision Cholesky with iterative refinement. Approximate $95\%$ confidence intervals are shown. Quarter precision is used on the coarsest discretisation level for the matrix factorisation, half precision on finer levels.
Figure 8: Cholesky + Iterative refinement MPML and double precision Cholesky MLMC memory access cost comparison. Total cost of the estimator per level is shown (cost of $1$ sample $\times$ number of samples). Quarter precision is used on the coarsest discretisation levels for the matrix factorisation, half precision on finer levels.

### v2

@@ Abstract
Mathematics Subject classifications. 65Y20, 65C05, 65C30, 65G20, 60-08.
@@ 3.2 Approximate FE solutions using inexact solvers
Due to the limitations of inexact linear solvers, the discrete solution $u_{h}$ is not obtained exactly in practice. Instead, we compute an approximation $\widehat{u}_{h}$. The aim of this section is to estimate the error $\|u_{h}(\cdot,\omega)-\widehat{u}_{h}(\cdot,\omega)\|_{H_{0}^{1}(D)}$ by means of the residual of the solution of the FE system $Ax=b$ and investigate how iterative refinement from Section 2.3 can be employed to solve this system. Let us first introduce some notation.
@@ 4.1 Standard multilevel Monte Carlo
Suppose $Q:\Omega\rightarrow\mathbb{R}$ is a random variable and we are interested in computing its mean $\mathbb{E}[Q]$. Assume that $Q$ cannot be evaluated sample-wise, but a sequence of models $Q_{l},l\in\{0,\ldots,N\}$, approximating $Q$ with increasing accuracy, is available. We define the following auxiliary Monte Carlo (MC) estimators:
@@
will then be referred to as the MLMC estimator for $E[Q]$. Although not necessary, the estimators $\widehat{Y}_{l}$ are, in this work, assumed to be independent. In this definition, it is assumed that the models can be evaluated exactly, i.e., we can compute $Q_{l}(\omega)$. In the following we sometimes abuse notation and denote by $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ an estimate computed using the estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$.
@@
It can be shown that three complexity regimes can be distinguished based on the values of $\beta$ and $\gamma$. The optimal performance of the MLMC algorithm is achieved when $\beta>\gamma$, i.e., when variance decays faster than cost increases. The full complexity analysis can be found in [cliffe2011multilevel, Theorem 1].
@@ 4.3 Application to the elliptic PDE problem
Let $G:H_{0}^{1}(D)\rightarrow\mathbb{R}$ be a bounded linear functional and consider Problem 3.1, an AVP with random data and solution $u(\cdot,\omega)\in H_{0}^{1}(D)$ for a.e. $\omega\in\Omega$. We consider the problem of estimating the quantity of interest (QoI) defined as the expected value of the random variable $Q:\Omega\rightarrow\mathbb{R}$ given by $\omega\mapsto G(u(\cdot,\omega))$.
@@
Let $m>1$, and let $h_{0},h_{1},\ldots$ be discretisation parameters satisfying $h_{0}>0$ and $mh_{l}=h_{l-1}$. Let $\widehat{u}_{h_{l}}$ be the discrete solution of the AVP with random data (Problem 3.1) computed effectively to precision $\delta_{l}$ on mesh level $l$, and assume that there are $k_{1},k_{2}>1$ such that $k_{1}\delta_{l}\leq\delta_{l-1}\leq k_{2}\delta_{l}$ for all $l\geq 1$. Then
@@ 4.4 Adaptive MPML algorithm
for a fixed constant $k_{p}\in(0,1)$. To balance the terms in the error estimate, it is sufficient to choose $\delta_{l},\;l=0,\ldots,L$ such that
@@
0:  $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}},\{N_{l}\}$
@@
7:     Update estimates for $N_{l}$ as $N_{l}\coloneqq\sqrt{\frac{V_{l}}{C_{l}}}\frac{2}{\text{TOL}^{2}}\sum_{k=0}^{L}\sqrt{V_{k}C_{k}}$
8:     if $|\widehat{Y}_{L}|>\frac{rm^{\alpha}-1}{\sqrt{2}}\text{TOL}$ then
@@
12:     if $|\widehat{Y}_{L}|\leq\frac{rm^{\alpha}-1}{\sqrt{2}}\text{TOL}$ and $\sum_{l=0}^{L}s^{2}_{l}/N_{l}\leq\text{TOL}^{2}/2$ then
@@ 5 Numerical results
We provide numerical experiments demonstrating the potential cost savings of the adaptive MPML algorithm (Algorithm 4.1) over a standard adaptive MLMC algorithm. The adaptive MPML algorithm preserves the overall computational accuracy. As a suitable model, we use an elliptic PDE with lognormal random coefficients with both a direct and an iterative linear solver. All numerical experiments are implemented in Python. The codes are available at https://github.com/josef-martinek/mpml.
@@ 00 $hhss$
Table 1: Left: Required effective precision on each level (in terms of relative residual), determined by the adaptive MPML algorithm (Algorithm 4.1) for different values of the finest level $L$.  Right: Choice of precision values $(\varepsilon_{f},\varepsilon_{7},\varepsilon,\varepsilon_{r})$ for iterative refinement on each level $l$. The factorisation is carried out in half precision on level 00.
As the quantity of interest we choose
@@
Due to the fact that $\omega_{j}\sim N(0,\sigma^{2})$, assumptions (4) and (5) are not satisfied. By choosing, e.g., $\omega_{j}\sim\text{Uni}(0,c)$, one could ensure that both assumptions are satisfied, but we want to test the developed methods in a more realistic setting. The analysis in Section 3.2 could easily be extended to this setting, using the analysis for MLMC presented in [teckentrup2013further], but we did not present this here to avoid unnecessary technicalities.
@@
In the first example of this section, we choose the data in (27) as follows. The right-hand side satisfies $f\equiv 1$ on $D$ and the parameters in the coefficient function are chosen as $s=4$, $q=2$, and $\sigma=2$. To solve this problem numerically, we discretise (27) using the FEM as described in Section 3 with simplical elements implemented in FEniCSx [baratta_2023_10447666]. In this experiment, the discretisation parameter on the coarsest mesh is $h_{0}=1/8$ and the mesh is refined on each level by a factor $m=2$.
@@
Recall that in all our experiments we refer to standard MLMC as Algorithm 4.1 without step 2, where the computations in step 4 are carried out with an a-priori given, sufficiently high accuracy. In each experiment we always describe precisely what this accuracy is. For MLMC, the parameters $\alpha$ and $\beta$ in the MLMC complexity theorem [cliffe2011multilevel, Theorem 1] (see also (10)) can be chosen to be $\beta=4$, $\alpha=2$ for the random field in (28). To set up the MPML algorithm (Algorithm 4.1), we further need the parameters $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, and $\beta_{2}$ from the MPML complexity theorem (Theorem 4.1). As shown in Lemma 4.3, we have $\alpha_{1}=\alpha=2$, $\beta_{1}=\beta=4$, $\alpha_{2}=1$, and $\beta_{2}=2$. In both adaptive algorithms, the underlying linear systems are solved using the PETSc implementation of MINRES [petsc_4_py] with a stopping criterion given by the relative residual norm. The tolerances for the stopping criterion in the MPML algorithm are given in Table 1 (left). The standard MLMC algorithm uses a fixed tolerance specified below for each experiment and all computations are carried out in double precision without iterative refinement. The cost of solving the linear system is estimated in terms of the number of floating point operations (FLOPs) performed by MINRES. To count the FLOPs we use the PETSc GetFlops() function.
@@
Figure 1: Estimated MSE with approximate $95\%$ confidence intervals for the MPMC estimator when compared to the reference MLMC estimator for various target tolerances. As their linear solver, both estimators use MINRES; MPML uses an adaptively chosen stopping criterion.
@@
Figure 2: Total cost gain in terms of FLOPs for various tolerances using adaptive MPML compared to the standard adaptive MLMC estimator. As their linear solver, both estimators use MINRES; MPML uses an adaptively chosen stopping criterion.
@@
To determine the effective precision $\delta_{l}$ in the MPML adaptive algorithm, we use formula (24) with the constant $k_{p}=0.05$. As discussed below and depicted in Figure 4, the MPML algorithm is not sensitive with respect to the choice of $k_{p}$. For a multilevel estimator with a given number of levels, the choice of $k_{p}=0.05$ then determines uniquely the effective precisions $\delta_{l}$ on all levels according to (24). The values of $\delta_{l}$ for different estimators can be found in Table 1 (left).
@@
We now aim to verify the accuracy of the MPML algorithm. We execute $1000$ runs of the adaptive MPML algorithm (Algorithm 4.1) for each of the four values of the MSE tolerance $\text{TOL}^{2}$, namely $\text{TOL}^{2}=8,4,2,1\times 10^{-6}$. For the same tolerances we perform $1000$ runs of the standard MLMC algorithm. The standard MLMC algorithm uses a fixed tolerance of $10^{-6}$ for the MINRES stopping criterion. Figure 1 shows the MSE of the estimates obtained by both algorithms averaged over the number of runs with approximate $95\%$ confidence intervals. To estimate the MSE of both algorithms a reference value of the QoI is computed by the standard MLMC algorithm with tolerance $\text{TOL}^{2}=2\times 10^{-8}$ and a direct, double precision linear solver. Up to statistical errors, both estimators have a similar MSE. However, they do differ slightly, due to the discrete choices in the two adaptive procedures, e.g., of the total number of levels $L$.
@@
In Figure 2, we demonstrate the cost gains obtained by MPML (Algorithm 4.1) compared to standard MLMC, again with a fixed tolerance of $10^{-6}$ in the MINRES stopping criterion. We plot the average cost in terms of FLOPs for each tolerance and observe that it is consistently reduced by a factor of $\approx 1.5$. As demonstrated in Figure 1, this cost gain comes with no statistically significant loss of accuracy. It results from the fact that the effective precision $\delta_{l}$ for solving the linear systems on the coarser levels is chosen significantly larger than $10^{-6}$ in MPML, see Table 1 (left).
@@
At the beginning of the cost analysis in Section 4.5 we made the assumption that the numbers of samples on each level are chosen approximately the same within the MPML and the MLMC adaptive procedures. We verify numerically that this is true. Figure 3 shows the average number of samples obtained by MPML (Algorithm 4.1) and standard MLMC on each level for the tolerance $\text{TOL}^{2}=10^{-6}$. We observe that the number of samples are equal within statistical error given by the approximate $95\%$ confidence intervals. For this experiment, the standard MLMC algorithm uses a fixed tolerance of $10^{-10}$ in the MINRES stopping criterion. Note that Level 4 is not reached in all runs and we need on average only one sample on Level 3.
@@
Figure 3: Average number of samples on each level for adaptive MLMC and adaptive MPML for $\text{\text{TOL}}^{2}=10^{-6}$ (with approximate $95\%$ confidence intervals).
@@
Figure 4: MSE of the MPML algorithm for various values of $k_{p}$ compared to the reference MSE of MLMC (with approximate $95\%$ confidence intervals).
@@
In Section 4.4 we claimed that the MPML algorithm is not sensitive to the a-priori chosen constant $k_{p}$ used to determine the required computational accuracy. Figure 4 shows the average MSE of the MPML algorithm (Algorithm 4.1) for the MSE tolerance $\text{TOL}^{2}=2\times 10^{-6}$ with approximate $95\%$ confidence intervals. The MSE is estimated using $1000$ runs of the algorithm for the values $k_{p}=0.05,0.1,0.2,0.4$. The estimated MSE of the standard MLMC algorithm is shown for comparison with a dashed line. It is estimated using $1000$ runs of standard MLMC with a tolerance of $10^{-10}$ in the MINRES stopping criterion. The choice of the constant $k_{p}$ seems to have no significant impact on the overall accuracy, provided the constant is chosen sufficiently small. For our suggested choice of $k_{p}=0.05$ the computational error in our model problem is bounded safely.
5.2 Elliptic PDE with lognormal coefficients – direct solver
@@
In this section, we again solve (27) with the same input data, i.e., $f\equiv 1$, $s=4$, $q=2$, and $\sigma=2$. The coarsest mesh size is again $h_{0}=1/8$. However, here the underlying linear system (with symmetric positive definite matrix) is solved using a double precision Cholesky factorisation from PETSc in the standard MLMC algorithm and a low precision Cholesky factorisation with iterative refinement in our MPML algorithm. Recall that throughout our experiments by standard MLMC we mean Algorithm 4.1 without step 2, where the computations in step 4 are (here) carried out with double precision. We use our own implementation of iterative refinement and of the low precision Cholesky factorisation. To carry out computations in half, single, and double precision, we use the numerical types of NumPy, namely float16(), float32(), and float64() for half, single, and double precision, respectively. Apart from the linear solver, all calculations are carried out in double precision.
@@
The iterative refinement (Algorithm 2.1) is set up as follows. As in the previous section, the actual, numerical values in the stopping criterion are given by the required effective precisions $\delta_{l}$ on each level, specified by formula (24). To this end, we also choose again $\alpha_{1}=\alpha=2$, $\beta_{1}=\beta=4$, $\alpha_{2}=1$, and $\beta_{2}=2$. Table 1 (left) shows the resulting values of the effective precision $\delta_{l}$. Table 1 (right) shows the exact setting of the other precisions used in the iterative refinement algorithm. The algorithm contains $3$ precisions, i.e., $\varepsilon_{f}$, $\varepsilon$, and $\varepsilon_{r}$, each taking on one of the values quarter ($q$), half ($h$), single ($s$), or double ($d$). To simplify the notation, we describe the exact setting of the iterative refinement schematically as an ordered quadruple, e.g., $(\varepsilon_{f},\varepsilon_{8},\varepsilon,\varepsilon_{r})=(hhss)$, where $\varepsilon_{8}$ is the precision chosen at step 8 of Algorithm 2.1.
In the cost model (10) we choose $\gamma=2$. In 2D, this corresponds to the case when the linear system is solved in linear complexity, since the number of unknowns grows quadratically with the mesh size.
@@
Figure 5: Comparison of estimated MSE for MPML and MLMC using the same number of samples and the same random seeds (with approximate $95\%$ confidence intervals). MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement.
@@
Figure 6: Comparison of total costs of the estimators per level in terms of memory access for MPML and MLMC. MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement.
@@
We start again by determining the accuracy of the MPML estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$, using the setting as described above, and we compare it to the accuracy of the standard MLMC estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ with double precision Cholesky as the linear solver on all levels. We perform $1000$ runs of the MPML estimator and of the MLMC estimator for each of the MSE tolerances $\text{TOL}^{2}=8,4,2,1\times 10^{-6}$. To eliminate randomness from estimating the MSE and to get a better idea of how big the computational error actually is, we use the same numbers of samples in the MPML and MLMC estimators and the same seeds in the random number generators. The number of samples $\{N_{l}\}$ we use for both estimators is determined by $1000$ runs of the standard adaptive MLMC algorithm. Figure 5 shows the resulting average mean squared errors of the MPML and MLMC estimators. The difference in the MSEs of both estimators is very small; in fact the relative error of the MPML MSE with respect to the MLMC MSE is less than $0.01\%$. Recall that the MPML estimator uses half and single precision for the matrix factorisation on all levels; see Table 1 (right). This promises significant memory savings when implemented efficiently on an architecture where half precision computations are supported.
@@
We continue by estimating the cost gains using MPML estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ with low precision Cholesky factorisation and iterative refinement compared to standard MLMC estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ with double precision Cholesky. Again, both estimators use the same numbers of samples $\{N_{l}\}$. Since memory references are by far the most costly part on modern computing architectures, both in terms of time and energy cost (cf. the discussion Section 2.4), we use memory access as a simple cost model, in terms of total number of bits loaded into main memory. Thus, accessing a half precision floating point number is 2×2$\times$2 × cheaper than accessing a single precision number and 4×4$\times$4 × cheaper than accessing a double precision number. The simulated cost gain can then be estimated by counting the number of entries in all vectors and the number of non-zeros in all sparse factorisations of the matrices computed and stored in the process of solving the linear systems during the iterative refinement process. Figure 6 shows the total cost gain per level (cost per sample $\times$ number of samples) for $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ compared to $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$. On all levels we observe a simulated memory gain of $\approx 2$. The total memory gain (sum over all levels) is $\approx 2$ as well.
@@
Figure 7: Comparison of estimated MSE for MPML and MLMC using the same number of samples and the same random seeds (with approximate $95\%$ confidence intervals). MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement (here with quarter precision on level 00).
Figure 8: Comparison of total costs of the estimators per level in terms of memory access for MPML and MLMC. MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement (here with quarter precision on level 00).